# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 3.2 Baseline simulation and statistics
We run a “short DNS” of length $T_{\text{short}}=4\times 10^{3}$ days 11$\approx$ 11≈ 11 years (after a 500-day spinup) to supply the pool of initially un-perturbed (“ancestral”) events. Then, to provide “ground truth” statistics, we run a control simulation, or “long DNS”, of duration $T_{\text{long}}=16\times 10^{3}\text{ days}\approx 44$ years, which is $O(1600)$ eddy turnover times and $O(160)$ jet meandering times (see Fig. 3 caption for timescale definitions). However, in estimating climatological statistics, we take advantage of statistical zonal symmetry by concatenating the timeseries of all 64 longitudes, increasing the effective sample size by a factor of $\sim L/$(some typical correlation length). Conceptually, the short and long DNS are analogous to “training” and “validation” datasets in standard machine learning procedures, in the sense that we want to infer properties of the validation set using only information extracted from the training set (for example, by perturbing and re-simulating events seen in training). As we show below, simply counting events from the short DNS gives probability estimates that deterioriate below $\sim\frac{1}{32}$, which we aim to rectify with boosting.
@@
Fig. 2 shows representative snapshots of three dynamical fields in the upper layer from the long DNS: tracer concentration $c$, zonal velocity $u=U-\partial_{y}\psi$, and meridional velocity $v=\partial_{x}\psi$. Fig. 3 shows Hovmöller diagrams of zonal-mean anomalies of $c$ and $u$ (not $v$, since zonal-mean meridional velocity is zero), as well as their climatological means and standard deviations plotted alongside the topography. These are statistics of the grid-cell values, not zonal means, but depend only on latitude because so does topography. Two eastward jets are prominent in the snapshots Fig. 2(b) and in the zonal mean profile Fig. 3b.iii, with preferred latitudes of $\sim\frac{1}{4}L$ and $\sim\frac{3}{4}L$. The Hovmöller diagram gives a sense of characteristic timescales: jets tend to remain roughly stationary for stretches of 100$\sim$ 100∼ 100 days at a time before shifting, as seen by the group of closed contours of $\psi$ and associated dipole of $u$ centered at time $t=3400$. and persisting ±50$\pm$ 50± 50 days to either side. Within these stretches of quasi-stationarity, there are shorter undulations of duration 10$\sim$ 10∼ 10, which we identify as the eddy turnover timescale.
@@
Figure 2: Snapshots of the QG system configuration in the upper layer. Contours indicate the anomaly streamfunction $\psi$, which varies over a non-dimensional range ±18$\pm$ 18± 18, dashed contours indicating negative anomalies. Colors indicate (a) tracer concentration $c$, (b) zonal wind velocity $u=U-\partial_{y}\psi$, where $U=1$ is the basic background shear, and (c) meridional velocity $v=\partial_{x}\psi$. The timestamps increase from left to right, and come from the long DNS. The small square represents an example target region in which to sample extremes of the local tracer concentration, in this case centered at $x_{0}=\frac{1}{2}L,y_{0}=\frac{26}{64}L$ and extending ±$\pm\ell=\frac{2}{64}L$ in both meridional and zonal directions. This same region is the target used in the following results.
@@
Figure 3: Hovmöller diagrams of anomalies (departures from time-means) of zonal-mean concentration (a.i) and zonal-mean zonal wind (b.i). Contours indicate zonal-mean streamfunction anomaly (range ±10$\pm$ 10± 10, negatives values dashed). Column (ii) shows bottom topography, which directly affects the lower layer only, but indirectly sets the preferred jet positions in the upper layer as well. Columns (iii) and (iv) show the climatological means and standard deviations of the same (not zonally averaged) quantities. The Hovmöller diagrams give context to the snapshots of $u$ from Fig. 2b, which come from times (i) 3300, when the upper and lower jets are both shifted south; (ii) 3400, when the jets are unusually far apart; and (iii) 3500, when the jets are unusually close together. These intermittent, discrete shifts in jet location happen every 100$\sim$ 100∼ 100 days, which we call the “jet meandering timescale”. During a typical 100-day timespan of stationary jet, the fields shown oscillate roughly 10 times; hence we assign the eddy turnover timescale a nominal value of 10 days.
@@ 3.3 Target variable
We define the intensity function of interest $R(\mathbf{x})$ as the upper-level concentration, $c_{1}$ (henceforth, simply $c$), averaged over a small square box $[x_{0}-\ell,x_{0}+\ell]\times[y_{0}-\ell,y_{0}+\ell]$ of half-width $\ell=\frac{2}{64}$, and 23 evenly spaced latitudes $y_{0}\in\big{\{}\frac{10}{64},\frac{12}{64},\ldots,\frac{54}{64}\big{\}}L$, restricted to the central region to avoid boundary effects. The central longitude $x_{0}$ is fixed to $L/2$, but by zonal homogeneity any longitude would be statistically equivalent. We also repeated the analysis with double the box length, and found results to be qualitatively similar. We will mostly show results only for the smaller box size. The effect of spatial scale is worth considering in its own right with a wider range, which we postpone to future work.
@@
While the low-order moments capture ordinary behavior of intensities $R$, the intensity peaks—i.e., severities $R^{*}$, defined in Sect. 2—are better viewed from an extreme value theory perspective, and summarized with the peaks-over-threshold procedure (Coles,, 2001). We set a threshold $\mu$ as the $(\frac{1}{2})^{5}$th complementary quantile of $R$, also denoted $\mu[(\frac{1}{2})^{5}]$, i.e., the level whose exceedance probability is $q(\mu)=(\frac{1}{2})^{5}$. Severities $R^{*}$ are extracted as cluster maxima above $\mu$, with buffer times $A_{\text{max}}=40$ days and $B=20$ days. All cluster maxima from the long DNS are used as input data points to infer the parameters (scale $\sigma$, shape $\xi$) of a generalized Pareto distribution (GPD), using the maximum-likelihood routine of the Extremes.jl package (Jalbert et al.,, 2024):
@@
[…]
@@
  • The GPD scale parameter, $\sigma$ is anti-correlated with $\mu$. The constraint $R^{*}\leq 1$ can explain this, as a higher threshold leaves less room for an expansive tail. Mathematically, a GPD tail can be adequately described by two different choices of threshold $(\mu_{1},\mu_{2})$, and the two corresponding scale parameters will be related by $\sigma_{2}-\sigma_{1}=\xi(\mu_{2}-\mu_{1})$. Only the shape parameter, $\xi$, is invariant with respect to $\mu$. For an upper-bounded tail, $\xi<0$ (as verified in Fig. 4f), hence $\sigma$ and $\mu$ vary inversely.
@@
We implemented the boosting and estimation procedures for every latitude separately, but for illustration focus the in-depth analysis on $y_{0}=\frac{26}{64}L$ (the small boxes in Fig. 2), an interesting location where the (detrended) mean is low, the threshold $\mu[(\frac{1}{2})^{5}]$ is low, the GPD scale $\sigma$ is large, and the GPD shape slightly more negative than in surrounding regions. Fig. 5 displays the underlying probability distributions at $y_{0}=\frac{26}{64}L$, clarifying the relationship between intensities, severities, and GPD parameters. The full PDF of intensity, in (a), has a positive skew and sub-Gaussian tail. Black and red dashed curves are estimates obtained from the long and short DNS, respectively, and 90% confidence intervals are obtained by longitudinal translation. Specifically, the shaded intervals are the 5th-95th percentile ranges of intensities at the same $y_{0}$, but with $x_{0}$ shifted from its base location of $\frac{1}{2}L$ by $\frac{0}{64}L,\frac{1}{64}L,\frac{2}{64}L,\ldots,\frac{63}{64}L$. The dashed black curve is the mean of all 64 curves, which effectively inflates the long DNS’s timespan by a factor of 64, to $T_{\text{long}}=64\times 16\times 10^{3}=1.024\times 10^{6}$ (divided again by some correlation length, whose precise value is not important for us here because we don’t aim for sped-up estimation—only correct estimation). The discrepancy between short and long DNS is most pronounced in the upper tail, which in panel (b) is magnified and integrated from the top, giving the CCDF. Gray lines mark the threshold, $\mu=0.52$, and its CCDF value $\frac{1}{32}\approx 0.03$. Above this level, the short DNS becomes rapidly more uncertain (error bar widens), and severely underestimates probabilities smaller than 0.005$\sim$ 0.005∼ 0.005.
@@
Both short and long DNS diverge markedly from the GPD fit shown in gray in panel (b). This is where the distinction between intensity and severity comes into play: the GPD is fitted to peaks over the threshold $\mu$—i.e., severities—which have a different distribution (specifically, shifted upward) than that of all exceedances over $\mu$, which would include the clusters surrounding the peaks. Panel (c) confirms that the GPD fits severities $R^{*}$ much better than it fits intensities $R$. If the threshold were raised, the clusters would shrink, the sequence of peaks would form a Poisson process, and the CCDFs of $R$ and $R^{*}$ would converge. For computational economy and because non-asymptotic extremes are of interest for climate risk, we keep the threshold at $\mu[(\frac{1}{2})^{5}]$ and formally define our goal with boosting as correcting the distribution of severities—not intensities. Hence, our measure of success will be whether the short-DNS severity CCDF in Fig. 5c, when augmented by boosting, will become closer to the long-DNS severity CCDF. The improved accuracy will be measured by $\chi^{2}$ divergence (18), with bins given by $r_{k}=\mu[(\frac{1}{2})^{5+k}]$ for $k=0,1,\ldots,10$.
@@
Figure 5: Probability distributions of local tracer concentrations at latitude $y_{0}=\frac{26}{64}L$ and averaged over a box of half-width $\ell=\frac{2}{64}L$. (a) The full PDF of intensity $R$; (b) the CCDF (tail integral) of intensity $R$, restricted to $R>\mu[\frac{1}{2}]$ (the median); (c) the CCDF of the severity $R^{*}$ (peaks of $R$ over the threshold $\mu[(\frac{1}{2})^{5}]$). Black and red dashed lines represent estimates from long and short DNS, respectively, with shaded 90% confidence intervals obtained by repeating the inference 646464 times, once for each possible longitudinal rotation of the dataset. The gray line in (b,c) represents the GPD fit to $R^{*}$.
@@ 4.1 Stochastic inputs
We perturb the QG model with impulsive forcing, as described generically in Sect. 2 and more concretely here by instantiating on the QG model. The stochastic input $\omega$ lives in the complex plane $\mathbb{C}$, and the state-space perturbation $G(\omega)$ consists of a single Fourier mode to be added to $\psi$. We choose the mode based on linear stability analysis, which is more easily explained as a procedure than as a closed formula:
@@
    Decompose $\psi$ into a Fourier basis $\psi_{z}(x,y)=\sum_{k,\ell}\widehat{\psi}_{z}(k,\ell)e^{i(kx+\ell y)}$, and write the linearized dynamics (about a state of rest, $\psi=0$) into the abstract form
@@
    Calculate the eigenvalues and eigenvectors $\{(\lambda^{(m)}(k,\ell),\widehat{\varphi}^{(m)}(k,\ell)):m=1,2\}$ of the Jacobian matrix $C^{-1}(k,\ell)D(k,\ell)$, ordered by stability: $\text{Re}\{{\lambda^{(1)}}\}\geq\text{Re}\{{\lambda^{(2)}\}}$, and select $(k^{*},\ell^{*})=\text{argmax}_{k,\ell}\{\text{Re}\{{\lambda^{(1)}(k,\ell)}\}$, i.e., the linearly most unstable mode from a rest state. Restrict the optimization to $(k,\ell)$ both nonnegative, and not both zero.
@@
    which can have pointwise magnitudes up to $2|\omega|$. In the QG model, the mode we identify is $(k^{*},\ell^{*})=(4,0)$, and $G(\omega)$ is plotted in Fig. 6c for three different example $\omega$s, which correspond to the points labeled 1,2,3 in panel (a). All share the same inter-layer relative phase and magnitude, as these are properties of $k^{*},\ell^{*}$, and $\widehat{\varphi}_{z}^{(1)}(k^{*},\ell^{*})$, but differ in absolute phase and magnitude. Note that points 2 and 3 are approximately diametrically opposed, and hence spatially $\sim 180^{\circ}$ out of phase, whereas point 1 is approximately one-quarter revolution away and spatially $\sim 90^{\circ}$ out of phase with both 2 and 3. Points (2, 3) are (closest to, farthest from) the center of the circle, and hence have the (smallest, largest)-magnitude spatial perturbations.
@@
The steps above completely specify $G(\omega)$, a linear map from $\mathbb{C}$ to functions of ($x,y,z$), which can be easily computed offline before running any ensembles. One could argue for two obvious refinements of this choice: (1) accounting for the non-zero background state by linearizing the quadratic form $J(q,\psi)$ and including that in the calculation of $D(k,\ell)$; and (2) accounting for finite time horizons by using the leading singular vector of the linear propagator, i.e., the initial infinitesimal error whose magnitude amplifies the most over a given time horizon (Farrell and Ioannou, 1996a, ; Farrell and Ioannou, 1996b, ). We demur on these suggestions, choosing to focus attention on the less-studied optimization of the advance split time given a fixed perturbation shape. There are several reasons that singular vectors may not be suitable for our goals. First, it is easier to compare different initial conditions, different advance split times, and even different topographies (which we don’t do here) when they are all subject to precisely the same perturbation. Second, as our results will demonstrate, the COAST tends to lie beyond the time range where linearized error dynamics are appropriate, which is natural because we aim for finite-amplitude boosts in extreme event amplitudes. Third, singular vectors are typically designed to optimize global errors, which might not be as relevant for local extremes. Fourth, such highly specialized perturbation shapes might not be accessible in a generic GCMs. Nonetheless, sensitivity analysis with respect to perturbation shape leads the agenda for follow-up work.
Having fixed a subspace $\Omega=\mathbb{C}$ for perturbations $\omega$, we need to specify an input distribution $p(\omega)$ over that space. We design the PDF for $\omega$ as a radially symmetric, smooth, compactly supported “bump function” parameterized by two scales: $W$ which is the maximum permissible magnitude of $\omega$, and $s$ which sets the typical perturbation strength:
@@
When $s\ll W$, $p$ is approximately a bivariate Gaussian density with diagonal covariance $s^{2}I$. When $s\gtrsim W$, $p$ is approximately uniform over the $W$-disc $\{\omega:|\omega|\leq W\}$, with rapid (but mathematically smooth) transition to 0 on the boundary. We fix $W=0.3$, limiting the maximum possible perturbation amplitude to $|\delta\psi|\leq 0.6$ (a characteristic streamfunction amplitude is $|\psi|\sim 10$). We include $s$ as a parameter to vary because there is no established principle to set the magnitude of impulses for the purpose of rare event sampling. In contrast, numerical weather prediction has an established (if heuristic) practice of tuning noise amplitude to match ensemble spread with model error (e.g., Berner et al.,, 2015). Optimizing for climatological accuracy is a different, murkier goal calling for less prejudice with regard to perturbation magnitude. We therefore vary $s$ widely from 0.060.060.06 to 0.90.90.9 in increments of 0.060.060.06 for 15 total values. $s$ is the impulsive-forcing analogue to the continuous-forcing amplitude that we called $F_{4}$ in Finkel and O’Gorman, (2024), which strongly influenced the perturbation growth rate and therefore the optimal advance split time.
Fig. 6(a,b) depicts $p(\omega;s,W)$ in two ways: (a) two-dimensional level sets of the unnormalized density (37) logarithmically spaced from $e^{-4}$ to $e^{-0.01}$, each value of $s$ occupying one of 15 sectors of the circle; and (b) one-dimensional transects across $p(\omega;s,W)$ fixing $\text{Re}\{\omega\}=0$. To save the labor of drawing Monte Carlo samples from $p(\omega;s,W)$ separately and simulating the perturbed children for each value of $s$, we compute the MoCTail and PoPTail estimators using numerical quadrature over the $W$-disc using a single set of samples drawn by quasi-Monte Carlo (QMC), and displayed as black dots in 6a. QMC is a general strategy which places samples deterministically across the input space in a way that mimics properties of randomness, but with lower discrepancy (fewer clumps and patches), thereby aiming to reduce variance in estimated statistics (Leobacher and Pillichshammer,, 2014). We specifically use the LatticeRuleSampler from the QuasiMonteCarlo.jl Julia library (Rackauckas,, 2023) to distribute points $\{(U_{m},V_{m})\}_{m=1}^{M}$ quasi-uniformly on the unit square $[0,1]^{2}$, and transform them to the $W$-disc with the formula
@@
which is the fraction of the $W$-disc between the radii $r_{1}$ and $r_{2}$. The phase $2\pi V$ is clearly $\mathcal{U}([0,2\pi])$. If $U$ and $V$ were independent random variables, we would immediately conclude $\omega$ is uniformly distributed over the $W$-disc; in QMC they are not independent, but the conclusion still holds true (Leobacher and Pillichshammer,, 2014). In all experiments to follow, $M=21$, corresponding to the 21 points plotted in Fig. 6a. While other sampling rules are possible, the LatticeRuleSampler enjoys a distinct advantage of being extensible: sampling 12 points at first and later deciding to add 9 more gives the same result as sampling 21 in one batch.

### v2

@@ 3.2 Baseline simulation and statistics
We run a “short DNS” of length $T_{\text{short}}=4\times 10^{3}$ days $\approx 11$ years (after a 500-day spinup) to supply the pool of initially un-perturbed (“ancestral”) events. Then, to provide “ground truth” statistics, we run a control simulation, or “long DNS”, of duration $T_{\text{long}}=16\times 10^{3}\text{ days}\approx 44$ years, which is $O(1600)$ eddy turnover times and $O(160)$ jet meandering times (see Fig. 3 caption for timescale definitions). However, in estimating climatological statistics from the long DNS, we take advantage of statistical zonal symmetry by concatenating the timeseries of all 64 longitudes, increasing the effective sample size by a factor of $\sim L/$(some typical correlation length). Conceptually, the short and long DNS are analogous to “training” and “validation” datasets in standard machine learning procedures, in the sense that we want to infer properties of the validation set using only information extracted from the training set (for example, by perturbing and re-simulating events seen in training). As we show below, simply counting events from the short DNS gives probability estimates that deterioriate at high levels of severity, which we aim to rectify with boosting.
@@
Fig. 2 shows representative snapshots of three dynamical fields in the upper layer from the long DNS: tracer concentration $c$, zonal velocity $u=U-\partial_{y}\psi$, and meridional velocity $v=\partial_{x}\psi$. Fig. 3 shows Hovmöller diagrams of zonal-mean anomalies of $c$ and $u$ (not $v$, since zonal-mean meridional velocity is zero), as well as their climatological means and standard deviations plotted alongside the topography. These are statistics of the grid-cell values, not zonal means, but depend only on latitude because so does topography. Two eastward jets are prominent in the snapshots Fig. 2(b) and in the zonal mean profile Fig. 3b.iii, with preferred latitudes of $\sim\frac{1}{4}L$ and $\sim\frac{3}{4}L$. The Hovmöller diagram gives a sense of characteristic timescales: jets tend to remain roughly stationary for stretches of $\sim 100$ days at a time before shifting, as seen by the group of closed contours of $\psi$ and associated dipole of $u$ centered at time $t=3400$. and persisting $\pm 50$ days to either side. Within these stretches of quasi-stationarity, there are shorter undulations of duration $\sim 10$, which we identify as the eddy turnover timescale.
@@
Figure 2: Snapshots of the QG system configuration in the upper layer. Contours indicate the anomaly streamfunction $\psi$, which varies over a non-dimensional range of approximately $\pm 18$, dashed contours indicating negative anomalies. Colors indicate (a) tracer concentration $c$, (b) zonal wind velocity $u=U-\partial_{y}\psi$, where $U=1$ is the basic background shear, and (c) meridional velocity $v=\partial_{x}\psi$. The timestamps increase from left to right, and come from the long DNS. The small square represents an example target region in which to sample extremes of the local tracer concentration, in this case centered at $x_{0}=\frac{1}{2}L,y_{0}=\frac{26}{64}L$ and extending $\pm\ell=\frac{2}{64}L$ in both meridional and zonal directions. This same region is the target used in the following results, and we consistently refer to the domain coordinates in fractions of 64 across all figures.
@@
Figure 3: Hovmöller diagrams of anomalies (departures from time-means) of zonal-mean concentration (a.i) and zonal-mean zonal wind (b.i). Contours indicate zonal-mean streamfunction anomaly (range $\pm 10$, negatives values dashed). Column (ii) shows bottom topography, which directly affects the lower layer only, but indirectly sets the preferred jet positions in the upper layer as well. For the same quantities, column (iii) shows the zonal and time mean and column (iv) shows the zonal mean of the temporal standard deviation. The Hovmöller diagrams give context to the snapshots of $u$ from Fig. 2b, which come from times (i) 3300, when the upper and lower jets are both shifted south; (ii) 3400, when the jets are unusually far apart; and (iii) 3500, when the jets are unusually close together. These intermittent, discrete shifts in jet location happen every $\sim 100$ days, which we call the “jet meandering timescale”. During a typical 100-day timespan of stationary jet, the fields shown oscillate roughly 10 times; hence we assign the eddy turnover timescale a nominal value of 10 days.
@@ 3.3 Target variable
We define the intensity function of interest $R(\mathbf{x})$ as the upper-level concentration, $c_{1}$ (henceforth, simply $c$), averaged over a small square box $x_{0}-\ell,x_{0}+\ell]\times[y_{0}-\ell,y_{0}+\ell]$ of half-width $\ell=\frac{2}{64}$, and 23 evenly spaced latitudes $y_{0}\in\big\{\frac{10}{64},\frac{12}{64},\ldots,\frac{54}{64}\big\}L$, restricted to the central region to avoid boundary effects. The central longitude $x_{0}$ is fixed to $L/2$, but by zonal homogeneity any longitude would be statistically equivalent. We also repeated the analysis with double the box length, and found results to be qualitatively similar, so we only show results for the smaller box size. The effect of spatial scale is worth considering in its own right with a wider range, which we postpone to future work.
@@
While the low-order moments capture ordinary behavior of intensities $R$, the intensity peaks—i.e., severities $R^{*}$, defined in Sect. 2—are better viewed from an extreme value theory perspective, and summarized with the peaks-over-threshold procedure (Coles, 2001). We set a threshold $\mu$ as the $(\frac{1}{2})^{5}$th complementary quantile of $R$, also denoted $\mu[(\frac{1}{2})^{5}]$, i.e., the level whose exceedance probability is $q(\mu)=(\frac{1}{2})^{5}$. Severities $R^{*}$ are extracted as cluster maxima above $\mu$, with buffer times $A_{\text{max}}=40$ days and $B=20$ days. All cluster maxima from the long DNS are used as input data points to infer the parameters (scale $\sigma$, shape $\xi$) of a generalized Pareto distribution (GPD), using the maximum-likelihood routine of the Extremes.jl package (Jalbert et al., 2024):
@@
Figure 4: Summary statistics of latitude-dependent climatological tail distributions of local tracer concentrations, also called “intensities”, which are denoted $R$ and defined as the average concentration $c$ over a box $(x,y)\in(x_{0},y_{0})+[-\ell,\ell]^{2}$. $x_{0}=\frac{1}{2}L$ and $\ell=\frac{1}{32}L$ are fixed, while $y_{0}$ varies across the midlatitudes from $\frac{10}{64}L$ to $\frac{54}{64}L$. Panel (a) shows the lower-layer topography in this same range of middle latitudes, (b) shows the mean intensity $\langle R\rangle(y_{0})$, after subtracting a nominal trend of $\frac{y_{0}}{L}$ to reveal a finer-scale structure that resembles the underlying topography, and (c) shows the standard deviation of intensity $\sqrt{\langle R^{2}\rangle-\langle R\rangle^{2}}$. Dashed curves in (b) and (c) indicate the mean and standard deviation, respectively, of the concentration field $c$ without box-averaging. Panels (d,e,f) summarize the distribution of intensities $R^{*}$ via the parameters of the generalized Pareto distributions (GPD), inferred by the peaks-over-threshold fitting procedure (see section 3.3 for details). The threshold is set to the $(\frac{1}{2})^{5}$-complementary quantile, denoted $\mu[(\frac{1}{2})^{5}]$ and shown in (d) with linear trend removed. Panels (e, f) display the estimated (scale, shape) parameters ($\sigma,\xi$).
@@
  • The GPD scale parameter, $\sigma$, is anti-correlated with the (detrended) mean. The constraint $R^{*}\leq 1$ can explain this, as a higher distribution center leaves less room for an expansive tail. In addition, the threshold $\mu$ tracks approximately with the mean, and we can understand the anticorrelation mathematically through the non-uniqueness of GPD parameters: the same tail can be adequately described by two different choices of threshold $(\mu_{1},\mu_{2})$, and the two corresponding scale parameters will be related by $\sigma_{2}-\sigma_{1}=\xi(\mu_{2}-\mu_{1})$. Only the shape parameter, $\xi$, is invariant with respect to $\mu$. Seeing that $\xi<0$ varies only slightly with latitude, $\sigma$ and $\mu$ would vary inversely even if the tail itself were not changing.
@@
We implemented the boosting and estimation procedures for every latitude separately, but for illustration focus the in-depth analysis on $y_{0}=\frac{26}{64}L$ (the small boxes in Fig. 2), an interesting location where the (detrended) mean and threshold $\mu[(\frac{1}{2})^{5}]$ are both low, the GPD scale $\sigma$ is large, and the GPD shape slightly more negative than in surrounding regions. Fig. 5 displays the underlying probability distributions at $y_{0}=\frac{26}{64}L$ to show the nature of the tails of the distributions and also to help clarify the relationship between intensities, severities, and GPD parameters. The full PDF of intensity, in (a), has a positive skew and sub-Gaussian tail. Black and red solid curves are estimates obtained from the long and short DNS, respectively, and 90% confidence intervals are obtained by longitudinal translation. Specifically, the shaded intervals are the 5th-95th percentile ranges of intensities at the same $y_{0}$, but with $x_{0}$ shifted from its base location of $\frac{1}{2}L$ by $\frac{0}{64}L,\frac{1}{64}L,\frac{2}{64}L,\ldots,\frac{63}{64}L$. The dashed black curve is the mean of all 64 curves, our best available estimate of ground truth. The discrepancy between short and long DNS is most pronounced in the upper tail, which in panel (b) is magnified and integrated from the top, giving the CCDF. Gray lines mark the threshold, $\mu=0.52$, and its CCDF value $\frac{1}{32}\approx 0.03$. Above this level, the short DNS becomes rapidly more uncertain (error bar widens), and severely underestimates probabilities smaller than $\sim 0.005$.
@@
Both short and long DNS estimates diverge markedly from the GPD fit shown in gray in panel (b). This is where the distinction between intensity and severity comes into play: the GPD is fitted to peaks over the threshold $\mu$—i.e., severities—whose distribution differs (most notably in the upward direction) from that of all exceedances over $\mu$, which would include the clusters surrounding the peaks. Panel (c) confirms that the GPD fit is much more appropriate for severities $R^{*}$ than for intensities $R$, and thereby clarifies the distinction. If the threshold were raised, the clusters would shrink, the sequence of peaks would form a Poisson process, and the CCDFs of $R$ and $R^{*}$ would converge. For computational economy and because non-asymptotic extremes are of interest for climate risk, we keep the threshold at $\mu[(\frac{1}{2})^{5}]$ and formally define our goal with boosting as correcting the distribution of severities—not intensities. Hence, our measure of success will be whether the short-DNS severity CCDF in Fig. 5c, when augmented by boosting, will become closer to the long-DNS severity CCDF.
@@
Figure 5: Probability distributions of local tracer concentrations at latitude $y_{0}=\frac{26}{64}L$ and averaged over a box of half-width $\ell=\frac{2}{64}L$. (a) The full PDF of intensity $R$. (b) The CCDF (tail integral) of intensity $R$, restricted to $R>\mu[\frac{1}{2}]$. (c) Further zoomed-in CCDF of the severity $R^{*}$ (peaks of $R$ over $\mu[(\frac{1}{2})^{5}]$). In all three panels, solid black and red lines represent estimates from long and short DNS, respectively, with shaded 90% confidence intervals obtained by repeating the inference $64$ times, once for each possible longitudinal rotation of the dataset. Error bars become degenerate at levels experienced by $5\%$ of longitudes. Black dashed lines show the mean over all longitudinal rotations, our best estimate of ground truth. The gray line in (b,c) represents the GPD fit to $R^{*}$ with $\mu=0.52$, $\sigma=0.06$, and $\xi=-0.31$, and this is a much better fit to the severities in (c) which makes sense given they are defined in terms of peaks.
@@ 4.1 Stochastic inputs
We perturb the QG model with impulsive forcing, which we now specify as a concrete instantiation of the generic form in Sect. 2. The stochastic input $\omega$ lives in the complex plane $\mathbb{C}(=\Omega,\text{ the ``input space''})$, and the state-space perturbation $G(\omega)$ consists of a single Fourier mode to be added to $\psi$. We choose the mode based on linear stability analysis, which is more easily explained as a procedure than as a closed formula:
@@
    Decompose $\psi$ into a Fourier basis $\psi_{z}(x,y)=\sum_{k,\ell}\widehat{\psi}_{z}(k,\ell)e^{i(kx+\ell y)}$, and write the linearized dynamics (about the baroclinically unstable background state with vertical zonal wind shear and $\psi=0$, and neglecting topography) into the abstract form
@@
    Calculate the eigenvalues and eigenvectors $\{(\lambda^{(m)}(k,\ell),\widehat{\varphi}^{(m)}(k,\ell)):m=1,2\}$ of the Jacobian matrix $C^{-1}(k,\ell)D(k,\ell)$, ordered by stability: $\text{Re}\{{\lambda^{(1)}}\}\geq\text{Re}\{{\lambda^{(2)}\}}$, and select $(k^{*},\ell^{*})=\text{argmax}_{k,\ell}\{\text{Re}\{{\lambda^{(1)}(k,\ell)}\}$, i.e., the linearly most unstable mode from the background state. Restrict the optimization to $(k,\ell)$ both nonnegative, and not both zero.
@@
    $G(\omega)=\delta\psi_{z}(x,y)=\omega\widehat{\varphi}_{z}^{(1)}(k^{*},\ell^{*})e^{i(k^{*}x+\ell^{*}y)}+\text{c.c.},$ (39)
@@
The steps above completely specify $G(\omega)$, a linear map from $\mathbb{C}$ to functions of ($x,y,z$), which can be easily computed offline before running any ensembles. One could argue for two obvious refinements of this choice: (1) specializing the linearization to the actual initial state, not just the background state, by linearizing the quadratic form $J(q,\psi)$ and including that in the calculation of $D(k,\ell)$; and (2) accounting for finite time horizons by using the leading singular vector of the linear propagator, i.e., the initial infinitesimal error whose magnitude amplifies the most over a given time horizon (Farrell and Ioannou, 1996a, b). We stick to the simpler approach of the most unstable modes of the background shear, choosing to focus attention on the less-studied optimization of the advance split time given a fixed perturbation shape. There are several reasons that singular vectors may not be suitable for our goals. First, it is easier to compare different initial conditions, different advance split times, and even different topographies (which we don’t do here) when they are all subject to precisely the same perturbation. Second, as our results will demonstrate, the COAST tends to lie beyond the time range where linearized error dynamics are appropriate, which is natural because we aim for finite-amplitude boosts in extreme events. Third, singular vectors are typically designed to optimize global errors, which might not be as relevant for local extremes. Fourth, such highly specialized perturbation shapes might not be accessible in a generic GCMs. Nonetheless, sensitivity analysis with respect to perturbation shape leads the agenda for follow-up work.
Having fixed a subspace $\Omega=\mathbb{C}$ for perturbations $\omega$, we need to specify an input distribution $p^{\Omega}(\omega)$ over that space. We design the PDF for $\omega$ as a radially symmetric, smooth, “bump function” which has compact support in order to prevent perturbations so large as to induce oscillatory transients. The PDF is parameterized by two scales: $W$ which is the maximum permissible magnitude of $\omega$, and $s$ which sets the typical perturbation strength:
@@
When $s\ll W$, $p$ is approximately a bivariate Gaussian density with diagonal covariance $s^{2}I$. When $s\gtrsim W$, $p$ is approximately uniform over the $W$-disc $\{\omega:|\omega|\leq W\}$, with rapid (but mathematically smooth) transition to 0 on the boundary. We fix $W=0.3$, limiting the maximum possible perturbation amplitude to $|\delta\psi|\leq 2W=0.6$ (see text after Eq. (39)), which is small compared to the characteristic streamfunction amplitude of $|\psi|\sim 10$. We include $s$ as a parameter to vary because there is no established principle to set the magnitude of impulses for the purpose of rare event sampling. In contrast, numerical weather prediction has an established (if heuristic) practice of tuning noise amplitude to match ensemble spread with model error (e.g., Berner et al., 2015). Optimizing for climatological accuracy is a different, murkier goal calling for less prejudice with regard to perturbation magnitude. We therefore vary $s$ widely from $0.06$ to $0.9$ in increments of $0.06$ for 15 total values. $s$ is the impulsive-forcing analogue to the continuous-forcing amplitude that we called $F_{4}$ in Finkel and O’Gorman (2024), which strongly influenced the perturbation growth rate and therefore the optimal advance split time.
Fig. 6(a,b) depicts $p(\omega;s,W)$ in two ways: (a) two-dimensional level sets of the unnormalized density (40) logarithmically spaced from $e^{-4}$ to $e^{-0.01}$, each value of $s$ occupying one of 15 sectors of the circle; and (b) one-dimensional transects across $p(\omega;s,W)$ fixing $\text{Re}\{\omega\}=0$. To save the labor of drawing Monte Carlo samples from $p(\omega;s,W)$ separately and simulating the perturbed children for each value of $s$, we compute the MoCTail and PoPTail estimators using numerical quadrature over the $W$-disc using a single set of samples drawn by quasi-Monte Carlo (QMC), and displayed as black dots in 6a. QMC is a general strategy which places samples deterministically across the input space in a way that mimics properties of randomness, but with lower discrepancy (fewer clumps and patches), thereby aiming to reduce variance in estimated statistics (Leobacher and Pillichshammer, 2014). We specifically use the LatticeRuleSampler from the QuasiMonteCarlo.jl Julia library (Rackauckas, 2023) to distribute points $\{(U_{m},V_{m})\}_{m=1}^{M}$ quasi-uniformly on the unit square $0,1]^{2}$, and transform them to the $W$-disc with the formula
@@
which is the fraction of the $W$-disc between the radii $r_{1}$ and $r_{2}$. The phase $2\pi V$ is clearly $\mathcal{U}([0,2\pi])$. If $U$ and $V$ were independent random variables, we would immediately conclude $\omega$ is uniformly distributed over the $W$-disc; in QMC they are not independent, but the conclusion still holds true (Leobacher and Pillichshammer, 2014). In all experiments to follow, $M=21$, corresponding to the 21 points plotted in Fig. 6a. While other sampling rules are possible, the LatticeRuleSampler enjoys a distinct advantage of being extensible: sampling 12 points at first and later deciding to add 9 more gives the same result as sampling 21 in one batch.